# SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models

### v1

@@ Abstract.
The rapid advancement of AI technologies, particularly Large Language Models (LLMs), is establishing a new paradigm for Business Intelligence (BI). Despite the emergence of pioneering work in enhancing BI systems with LLMs, we have identified the following three issues when deployed in real industrial scenarios: interaction limitations, performance bottlenecks, and functionality deficiencies.
In this paper, we present SiriusBI, an end-to-end business intelligence system that is designed to address the three issues simultaneously. First, we propose an intelligent and application-oriented module called multi-round dialogue with querying, which aims to overcome the prevalent interaction limitations in current BI solutions. Next, to mitigate the performance bottlenecks caused by scenario migration, we introduce two SQL generation methods that strike a balance between accuracy and deployment costs. Finally, to tackle the practical challenges posed by functionality deficiencies, we develop an end-to-end workflow that covers the entire BI process, ensuring that SiriusBI delivers a robust and complete set of functionalities.
As an independent cloud service in Tencent’s data platform, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients. Experiments on real-world datasets and practical applications in industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI. Remarkably, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
@@
Jie Jiang,
Haining Xie,
Yu Shen,
Zihan Zhang,
Meng Lei,
Yifeng Zheng,
Yide Fang,
Chunyou Li,
Danqing Huang,
Wentao Zhang,
Yang Li,
Xiaofeng Yang,
Bin Cui,
Peng Chen
. PVLDB, 14(1): XXX-XXX, 2020.
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
Business Intelligence (BI) (Negash, 2004; Wieder and Ossimitz, 2015) is a crucial application scenario in the data field, comprising a comprehensive suite of methodologies, tools, and infrastructures designed to collect, integrate, analyze, and present an organization’s raw data to generate actionable insights for informed decision-making. The primary objective of BI is to generate actionable insights that facilitate informed decision-making. BI systems are extensively utilized across various sectors, including finance (Nuseir, 2021), environment (Hegger and Saraceni, 2024), and social media (Chen et al., 2024a; Rivera, 2024), which significantly enhances the decision-making process through the provision of real-time analytics and reporting capabilities.
@@
A typical BI system comprises several key components: a data management system that stores, processes, and aggregates vast amounts of data; analytics algorithms that transform this data into actionable insights; and visualization tools that present the information in intuitive and user-friendly formats. These components collaboratively function to streamline business processes, enhance customer experiences, reduce operational costs, and mitigate risks. Recent advancements in LLMs have sparked significant interest in ChatBI (Lian et al., 2024) — an integrated BI system enhanced by conversational AI. Concurrently, the demand for a fully integrated and efficient ChatBI solution is surging, driven by the need of a more intuitive and accessible mode of data interaction. This evolution promises to transform how users engage with data, making insights more readily available and actionable.
To meet the growing demand for big data analysis and decision-making in BI, the data community has proposed numerous effective approaches. While NL2SQL is the core task in most BI systems, traditional methods often rely on the schema-based approaches (Hristidis and Papakonstantinou, 2002; Luo et al., 2007; Zeng et al., 2016; Hristidis et al., 2003) or parsing-based approaches (Iyer et al., 2017; Wang et al., 2017; Li and Jagadish, 2014; Popescu et al., 2004) to generate SQL queries from natural language. Recently, there has been a growing trend in implementing NL2SQL powered by LLMs, utilizing either prompt engineering (Pourreza and Rafiei, 2023; Wang et al., 2023) or model training techniques (Li et al., 2024; Pourreza and Rafiei, 2024). However, these methods typically focus on Single-Round Dialogue (SRD) scenarios, which limits their effectiveness in real-world applications where user interactions are typically more complex. To address the limitations of NL2SQL in BI contexts, the NL2BI (Lian et al., 2024) task has been introduced. This task aims to generate SQL queries based on Multi-Round Dialogues (MRD), allowing for a comprehensive understanding of user intentions. However, when implementing existing work in industrial BI systems and applying it in real-world scenarios, we identified the following three challenges:
C1: Interaction Limitations. ChatBI aims to leverage MRD to assist users in data analysis, thereby lowering the barrier to data usage and enhancing overall data utilization efficiency. However, existing public solutions still face challenges in supporting MRD due to real-world interaction modes. On one hand, much of the research (Wang et al., 2023; Talaei et al., 2024; Li et al., 2024; Pourreza and Rafiei, 2024, 2023) in the community concentrates on enhancing the performance of the NL2SQL task. While this has advanced the field to some extent, the interaction mode remains restricted to SRD, which imposes strict constraints on how users express their needs. This limitation ultimately undermines the primary objective of ChatBI systems: to facilitate universally accessible data analysis. On the other hand, leading technology teams in industry have introduced ChatBI (Lian et al., 2024) that support MRD and offered feasible methods for implementing multi-round interactions. Nevertheless, these solutions also impose limitations on user interaction modes in practical applications, catering only to fixed question-answer formats. Furthermore, they tend to be less effective in BI scenarios that require deep domain expertise.
C2: Performance Bottlenecks. The NL2SQL task is crucial for BI because it directly impacts the accuracy of data retrieval during analysis. In practical industrial scenarios, existing NL2SQL methods often exhibit suboptimal performance when applied across diverse contexts. For example, a LLM fine-tuned on data specifically annotated for the financial domain performs poorly when applied to the advertising domain. The process of preparing and standardizing data for each specialized field, followed by fine-tuning, incurs significant costs, which are often impractical in industrial applications.
C3: Functionality Deficiencies. Traditional BI systems are typically comprehensive solutions for data analysis, encompassing multiple interrelated modules such as data storage, SQL generation, and data insights. However, existing LLM-based BI systems fall short in providing essential functionalities, including efficient intent understanding, adaptive domain-specific knowledge, and intelligent data insights. For example, Mac-SQL (Wang et al., 2023) and CHESS (Talaei et al., 2024) primarily focus on NL2SQL tasks, while Lian et al. (2024) propose using the Apache SuperSet(Apa, [n.d.]) tool directly after NL2SQL, neglecting the components of domain knowledge management module and deep data insight. Unlike traditional BI systems, users expect LLMs to enhance the entire ChatBI workflow, which drives a technological paradigm shift that improves both application efficiency and user experience. A purely NL2SQL-based approach or scenario-constrained MRD functionality is inadequate to meet the real-world needs of BI applications. This gap presents an opportunity to develop a more comprehensive and intelligent ChatBI system.
@@
To address the aforementioned challenges, we propose SiriusBI, which implements an end-to-end solution for BI scenarios. This system leverages the capabilities of LLMs to empower various submodules, thereby enhancing both efficiency and user experience in data analysis. Specifically, for the issue of interaction limitations (C1), we introduce the MRD-Q (Multi-Round Dialogue with Querying) module. Unlike the MRD functionality proposed by Lian et al. (2024), MRD-Q incorporates an intent clarification module that disambiguates user queries through follow-up questions. This approach enables the system to accurately identify the user’s true intent, even when the initial query is flawed or ambiguous, thereby facilitating precise responses, as shown in Figure 1. Furthermore, we employ a Retrieval-Augmented Generation (RAG) framework that leverages both business-domain knowledge and metadata, ensuring our system’s adaptability across various domains. Consequently, this reduces barriers for users and significantly enhances the overall user experience.
For the issue of performance bottlenecks (C2), we propose two implementation strategies in SQL generation. First, for domain-specific NL2SQL tasks, we introduce an efficient and automated process for constructing training data to improve the performance of LLMs. Second, for scenarios that require transferability, we present a two-step solution that integrates a Semantic Model with a SQL Model. This approach leverages domain-specific knowledge and metadata, ensuring high accuracy in SQL generation without the need of additional training. Furthermore, it reduces the costs associated with transferring between different scenarios and enhances the generalization of NL2SQL across various domains.
For the issue of functionality deficiencies (C3), SiriusBI provides a comprehensive end-to-end workflow. Unlike existing works that mainly focus on the NL2SQL task and the expansion of MRD modules, SiriusBI introduces two essential modules specifically designed for ChatBI : the Knowledge Management module and the Data Insights module. The knowledge management module is responsible for storing, processing, and utilizing domain-specific expertise and data to support the specialized knowledge requirements of various submodules. Meanwhile, the Data Insights module transcends traditional data visualization and report generation by providing critical key functionalities such as automatic data preparation, complex task planning, and the application of insight tools. These capabilities provide users with more comprehensive and in-depth support for their data analysis needs.
@@
We propose SiriusBI, an end-to-end LLM-based ChatBI system designed to provide better data analysis services and lower the usage threshold for users. To the best of our knowledge, SiriusBI is the first comprehensive end-to-end ChatBI system that simultaneously addresses the aforementioned three challenges. So far, SiriusBI has been applied across Tencent’s finance, advertising, and cloud sectors, providing services to dozens of enterprise clients.
To extend the interaction modes of ChatBI systems, we introduce MRD-Q, a dialogue system designed to ask clarifying questions when user queries are unclear or ambiguous. This improves the accuracy of intent understanding, lowers the barrier to user entry, and ultimately enhances the overall user experience.
To enhance the transferability of ChatBI systems across different industrial scenarios, we propose two implementation approaches for SQL generation: single-step generation and two-step generation. The two-step generation method significantly improves the transferability of ChatBI and provides a new paradigm for SQL generation in the industry.
@@
Extensive experiments on real-world datasets and industrial BI scenarios demonstrate the practicality and effectiveness of SiriusBI. Notably, when applied in industry, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 89% for Tencent Advertisement, and 91% for Tencent Cloud.
2. Preliminary
In this section, we introduce some key concepts including BI, ChatBI, NL2SQL, knowledge management and data insight.
2.1. BI and ChatBI
BI refers to the technologies, applications, and practices involved in the collection, integration, analysis, and presentation of business information. The primary goal of BI is to improve decision-making with business organizations. Traditional BI systems typically rely on data warehousing, data mining, and reporting tools, which enable organizations to transform raw data into meaningful insights.
ChatBI represent an evolution of traditional BI systems by integrating natural language processing (NLP) capabilities. This integration allows users to interact with BI tools using conversational interfaces. ChatBI systems aim to democratize data access by enabling non-technical users to query and analyze data through natural language queries. These systems leverage advanced NLP techniques to accurately understand user intent and generate relevant responses, thereby enhancing the overall user experience and allow organizations to make data-driven decisions more accessible and actionable.
2.2. NL2SQL Task
NL2SQL is a task in the field of natural language processing and database management that focuses on converting natural language queries into SQL queries.
In the specific task of SQL generation, we first receive the user’s query intent . Then, based on the relevant metrics and dimensions within the user’s intent, and in conjunction with the database schema information , and the SQL query result can be generated as,
where is the function that generates the SQL query result . The accuracy and efficiency of NL2SQL systems are critical for the effectiveness of ChatBI systems, as they directly impact the quality of the generated insights.

### v2

# SiriusBI: A Comprehensive LLM-Powered Solution for Data Analytics in Business Intelligence

@@ Abstract.
With the proliferation of Large Language Models (LLMs) in Business Intelligence (BI), existing solutions face critical challenges in industrial deployments: functionality deficiencies from legacy systems failing to meet evolving LLM-era user demands, interaction limitations from single-round SQL generation paradigms inadequate for multi-round clarification, and cost for domain adaptation arising from cross-domain methods migration.
We present SiriusBI, a practical LLM-powered BI system addressing the challenges of industrial deployments through three key innovations: (a) An end-to-end architecture integrating multi-module coordination to overcome functionality gaps in legacy systems; (b) A multi-round dialogue with querying mechanism, consisting of semantic completion, knowledge-guided clarification, and proactive querying processes, to resolve interaction constraints in SQL generation; (c) A data-conditioned SQL generation method selection strategy that supports both an efficient one-step Fine-Tuning approach and a two-step method leveraging Semantic Intermediate Representation for low-cost cross-domain applications. Experiments on both real-world datasets and public benchmarks demonstrate the effectiveness of SiriusBI. User studies further confirm that SiriusBI enhances both productivity and user experience.
As an independent service on Tencent’s data platform, SiriusBI is deployed across finance, advertising, and cloud sectors, serving dozens of enterprise clients. It achieves over 93% accuracy in SQL generation and reduces data analysts’ query time from minutes to seconds in real-world applications.
@@
PVLDB, 14(1): XXX-XXX, 2020.
@@
The source code, data, and/or other artifacts have been made available at https://github.com/TencentBigData-SiriusAI/SiriusBI.
@@ 1. Introduction
Business Intelligence (BI) (Negash, 2004; Wieder and Ossimitz, 2015) is a crucial application scenario in the data field, comprising a comprehensive suite of methodologies, tools, and infrastructures designed to collect, integrate, analyze, and present raw data from an organization to generate actionable insights for informed decision-making. BI systems are extensively used in various sectors, including finance (Nuseir, 2021), environment (Hegger and Saraceni, 2024), and social media (Chen et al., 2024a; Rivera, 2024), which significantly improves the decision-making process through the provision of real-time analytics and reporting capabilities (Wang et al., 2025; Li, 2024).
@@
A typical BI system comprises several key components: a data management module that stores, processes, and aggregates vast amounts of data; analytic algorithms that transform the data into actionable insights; and visualization tools that present the information in intuitive and user-friendly formats. Among these, data analytics plays a crucial role in providing decision-making support, directly determining the correctness and appropriateness of decisions. Recent advancements in LLMs (Lian et al., 2024; Ke et al., 2024; Zhang et al., 2024) have sparked significant interest in ChatBI — a new paradigm supported by natural language interfaces (Affolter et al., 2019). Concurrently, the demand for a fully integrated and efficient ChatBI solution is surging, driven by the need of a more intuitive and accessible mode of data interaction. This evolution promises to transform how users engage with data, making insights more available and actionable.
To meet the growing demand for big data analytics and decision-making in BI, the data community has proposed numerous effective approaches. However, when applying existing work in real-world BI scenarios, we identify the following three challenges:
C1: Functionality Deficiencies. While traditional business intelligence systems (Chaudhuri et al., 2011) integrate core components spanning data management, SQL generation, and insight discovery to form complete analytics pipelines, their reliance on heuristic rules and conventional AI/ML techniques limits generalization ability in dynamic scenarios. Although LLM-based methods have advanced task-specific performance, few offer comprehensive BI capabilities comparable to their traditional counterparts. For example, MAC-SQL (Wang et al., 2023b) and CHESS (Talaei et al., 2024) optimize NL2SQL accuracy but treat SQL execution as terminal outputs, neglecting downstream tasks like attribution analysis. While Lian et al. (Lian et al., 2024) extend their pipeline with Apache Superset for visualization, they fail to introduce knowledge bases to support dynamic grounding of domain-specific context, a critical requirement for real-world BI adaptation (Cheng et al., 2009). This functional fragmentation forces users to manually coordinate tools (e.g., SQL editors, dashboard platforms, knowledge retriever), which imposes significant cognitive load and reduces operational efficiency, as evidenced by industry report (Kurt Schlegel, 2023).
C2: Interaction Limitations. In the context of ChatBI, the NL2SQL task is becoming increasingly vital, as it facilitates seamless interaction between natural language queries and structured data retrieval, thereby enhancing the efficiency and accuracy of data analytics. The evolution of NL2SQL techniques reveals a critical architectural mismatch: while traditional methods (schema-based (Hristidis and Papakonstantinou, 2002; Luo et al., 2007; Zeng et al., 2016; Hristidis et al., 2003) or parsing-based (Iyer et al., 2017; Wang et al., 2017; Li and Jagadish, 2014; Popescu et al., 2004)) and modern LLM-driven approaches (prompt engineering (Pourreza and Rafiei, 2023; Wang et al., 2023b) or fine-tuning techniques (Li et al., 2024c; Pourreza and Rafiei, 2024)) predominantly optimize for Single-Round Dialogue (SRD) precision. This SRD-centric paradigm introduces a significant continuity gap in Multi-Round Dialogues (MRD): real-world BI workflows often require iterative investigation through successive queries, where later queries tend to omit previously provided contextual information, resulting in semantic ambiguity beyond the initial query. For instance, in the MRD NL2SQL task illustrated in Figure 1, the user issues three queries; notably, the second and third queries omit the time condition “2023” because it was specified in the first query. Single-round NL2SQL approaches demand nearly perfect input specificity, which explains its failure to generate correct SQL statements for the second and third queries in Figure 1. Worse still, due to the intricate nature (Liu et al., 2024) of MRD, few approach has been devoted to addressing this task. Lian et al.’s MRD solution (Lian et al., 2024) is the first attempt towards this task. Nevertheless, their solution is absent of user-guided clarification loops for intent resolution and domain-grounded dialogue act modeling. As demonstrated in Figure 1, the basic MRD approach still exhibit performance degradation beyond the first dialogue round.
C3: Cost for Domain Adaptation. Cross-domain deployment of NL2SQL models faces the challenge of cost surges in domain knowledge transfer, primarily caused by insufficient model generalization capability. Structural differences in database schema across domains (e.g., nested tables in finance vs. wide tables in advertising) necessitate repetitive model adaptation (Liu et al., 2024), while semantic gaps between industry-specific operators (e.g., financial window functions vs. e-commerce promotional rules) exacerbate logical deviations in SQL generation (Li et al., 2024a). Critically, domain knowledge transfer relies heavily on expert-annotated data, with manual annotation costs growing with domain complexity (Li et al., 2024b). Our real-world deployment statistics show that direct model migration leads to business logic errors in approximately two-thirds of generated SQL queries. Meanwhile, adapting models through traditional fine-tuning requires 5.5 person-days on average to label 200 seed queries within existing databases—forming critical bottlenecks for enterprise-level scalability.
To address the aforementioned challenges, we propose SiriusBI, which implements a comprehensive LLM-powered solution for ChatBI scenarios. This system leverages the capabilities of LLMs to empower various modules, thereby enhancing both the efficiency and user experience in data analytics. Specifically, for the issue of functionality deficiencies (C1), SiriusBI introduces an end-to-end integrated architecture that seamlessly orchestrate core modules including knowledge management, multi-round dialogue analysis, SQL generation, and data insight provision, thereby ensuring a closed-loop pipeline from natural language queries to final decision-making reports.
@@
For the issue of interaction limitations (C2), we introduce the MRD-Q (Multi-Round Dialogue with Querying) module. As a supplement to the basic multi-round dialogue analysis module proposed by Lian et al. (Lian et al., 2024), MRD-Q incorporates an intent querying module to clarify user queries through follow-up questions. This approach enables the system to accurately identify the user’s true intent, even when the initial query is incomplete or ambiguous, thus facilitating precise responses, as shown in Figure 1.
To enable economic domain adaptation (C3), SiriusBI introduces a strategy switching mechanism that dynamically selects between one-step and two-step SQL generation paradigms based on data conditions. This mechanism optimizes the trade-off between computational cost and performance, ensuring efficient adaptation to diverse business domains. To address structural and semantic disparities across domains, the one-step method employs a supervised fine-tuning (SFT) method enhanced by an automated data generation process pipeline. This pipeline minimizes manual annotation costs by leveraging SQL logs and LLM-based reverse engineering to synthesize high-quality training data, greatly reducing reliance on labor-intensive expert labeling. For cold-start scenarios, the two-step method eliminates SFT by combining domain knowledge retrieval with semantic intermediate representation (SIR), achieving robust performance across diverse industries without domain-specific tuning. The generalization ability of our system is further enhanced by a knowledge base with hybrid storage strategies, specially designed for BI application scenario.
@@
Practical ChatBI System: SiriusBI integrates four core functionalities — knowledge management, multi-round dialogue with querying, SQL generation, and data insight — into a unified system. It has been deployed across Tencent’s finance, advertising, and cloud businesses departments, serving dozens of enterprise clients, reducing data analysis latency from minutes to seconds.
New Dialogue Analysis Mechanism: We propose a new dialogue analysis mechanism named MRD-Q, consisting of semantic completion, knowledge-guided clarification, and proactive querying processes. It effectively addresses issues raised by ambiguous or incomplete queries in real-world scenarios.
Economic Domain Adaptation Strategy: We propose a dynamic strategy selection mechanism for SQL generation conditioned on data characteristics. To reduce annotation costs for fine-tuning, we set up an automated data synthesis pipeline. To enhance cross-domain SQL syntax accuracy and business logic consistency, we use Semantic Intermediate Representation (SIR) in prompt engineering.
Multi-Round NL2SQL Benchmark: We construct MRD-BIRD —a dataset containing 96 multi-round dialogue sequences (3–5 interaction rounds)— and open-source the data to advance NL2SQL research in multi-round dialogue contexts.
@@
Proven Effectiveness: Extensive experiments on both academic benchmarks and real-world datasets demonstrate the effectiveness of SiriusBI. Notably, when applied in industry, SiriusBI achieves remarkable accuracy rates of 97% in SQL generation for Tencent Finance, 93% for Tencent Advertisement, and 96% for Tencent Cloud. User studies further confirm that SiriusBI enhances both productivity ad user experience.
2. Related Work
2.1. Business Intelligence
The integration of AI and Machine Learning (ML) has transformed data-driven decision-making in BI (Rane et al., 2024). ML and Predictive Analytics enhance forecasting and pattern recognition in large datasets (Lee et al., 2022; Tamang et al., 2021). NLP and Text Analytics extract insights from unstructured data, improving customer service and decision-making (Kaur et al., 2019; Ittoo et al., 2016). The rise of the ChatBI paradigm, which has significantly advanced business intelligence and taken a major step towards data democratization (Linkon et al., 2024), is largely driven by the development of advanced AI models such as BERT (Devlin et al., 2019) and ChatGPT (team, 2022), which enable intelligent data interpretation and improved user interactions. Among related works, BI-REC (Meduri et al., 2021) is a pioneering system that recommends BI applications to help users achieve their data analysis goals. Similar to our method, BI-REC includes a step for predicting BI intents. However, BI-REC focuses on leveraging historical state information to predict possible future BI intents, whereas our approach aims to assess whether the current user intent is sufficiently clear and precise to generate an accurate SQL query. Lian et al. (Lian et al., 2024) are the first to recognize the importance of the NL2BI task in practical ChatBI applications and propose an efficient method to address it. We draw inspiration from their dialogue analysis technique to obtain semantically complete queries. The key differences between our approach and theirs are: (1) we leverage LLMs instead of smaller pre-trained models to analyze query completeness; (2) we incorporate domain-specific knowledge to help the LLM better understand user intentions; and (3) we enable active clarification questions directed at users when the LLM determines that the query requires further disambiguation. Lian et al.’s method does not consider the latter two aspects.
2.2. NL2SQL Methods
NL2SQL is a core task in ChatBI systems and can be broadly categorized into prompt-based and training-based methods (Zhou et al., 2024). Recently, Large Language Models (LLMs) like Codex (team, 2023a) and Claude 3 (Team, 2024) have shown great promise in NL2SQL, leveraging techniques such as prompt engineering (Giray, 2023), in-context learning (Dong et al., 2022), and chain-of-thought prompting (Wei et al., 2022). Prompt-based approaches not only offer new technical paradigms but also reduce costs in industrial settings. Notable examples include DIN-SQL (Pourreza and Rafiei, 2023), C3 (Dong et al., 2023), MAC-SQL (Wang et al., 2023b), and SQL-PaLM (Sun et al., 2023), which improve NL2SQL performance on models like ChatGPT and PaLM (Narang and Chowdhery, 2022).
Before LLMs, NL2SQL solutions mainly relied on training encoder-decoder sequence-to-sequence models (Cai et al., 2021; Cao et al., 2021; Choi et al., 2021; Gan et al., 2021; Wang et al., 2020a), requiring large amounts of database and SQL data. The advent of LLMs has introduced new training paradigms; for instance, Codes (Li et al., 2024c) proposes a training method tailored for NL2SQL that boosts LLM performance. However, some studies (Maamari et al., 2024; Pourreza et al., 2024; Caferoğlu and Ulusoy, 2024) fine-tune LLMs like ChatGPT (team, 2022) and Gemini (Team et al., 2023) to further improve results, though this often incurs high practical costs.
3. Method
3.1. Overview
The overall framework of SiriusBI is illustrated in Figure 2 and consists of four core modules:
(1) The Knowledge Management module consists of four types of data – database metadata, domain-specific knowledge, task-related materials (e.g., demonstrations used for in-context learning in LLMs), and historical conversations. These data types are stored in a hybrid architecture, ensuring efficient and scalable access to knowledge, supporting the seamless functionality of other modules.
(2) The Multi-Round Dialogue with Querying (MRD-Q) module is composed of two components: Dialogue Analysis and Intention Querying, which enable multi-turn user-oriented dialogue and knowledge-guided clarification to resolve semantic ambiguity in Conversational BI. Unlike existing SRD-centric approaches (Wang et al., 2023b; Pourreza and Rafiei, 2023), MRD-Q introduces follow-up question generation based on domain-specific knowledge, enabling context-aware SQL calibration across 3-5 dialogue turns.
(3) The SQL Generation module dynamically selects between one-step (fine-tuning based) and two-step (semantic intermediate representation based) processing paradigms according to business scenario data characteristics. This dual-mode approach efficiently transforms user intent into SQL queries with optimal cost-effectiveness. The one-step method streamlines the process through an automated data preparation pipeline, while the two-step method employs a knowledge-base-driven query rewriting strategy that eliminates model training requirements. This architectural design balances operational efficiency with implementation flexibility across diverse application contexts.
(4) To close the loop from SQL execution to decision support, SiriusBI integrates a Data Insight module that interprets query results through a multi-agent workflow based on the FunctionCall (Kim et al., 2024) and ReAct (Yao et al., 2023) mechanism as shown in Figure 3. Specifically, the Planner Agent analyzes user queries, SQL results, and domain knowledge to decompose tasks into subtasks and issue instructions; the Data Preparation Agent identifies missing data requirements and generates SQL queries to retrieve necessary information; the Tool Execution Agent selects specialized tools (Forecast/Diagnosis/Attribution Tools, implemented via reproduced existing works (Box et al., 2015; Hundman et al., 2018; Bhagwan et al., 2014; Li et al., 2019; Lundberg and Lee, 2017)) to process data, with final outputs consolidated by the Planner Agent for response generation.